# AM-RADIO: Agglomerative Vision Foundation Model Reduce All Domains Into One

### v4

@@ Abstract
Figure 1: PCA feature visualization of different models. Our proposed RADIO model can process any resolution and aspect ratio, and produces semantically rich dense encodings. Not only is the model capable of producing state-of-the-art feature representations at arbitrary resolutions, but the adaptor heads may be used to simulate other models at resolutions they don’t support, such as SAM at larger than 1024px.
@@ 4 Implementation Details
Multi-Resolution ViTs. Many of our student models use ViT [15] as the base vision architecture. Traditionally, ViTs use a learned position embedding for each input patch in an image, which in turn enforces that the model always operates at a constant resolution. We employ the Cropped Position Embedding (CPE) [30] augmentation with the number of positions being equal to $128^{2}$. The position embeddings are then randomly cropped and interpolated to match the number of input patches for the student model. Even when training with CLIP+DINOv2 at 224 resolution, we found that this technique results in a negligible drop (Table 5) in summary metrics, but improved semantic segmentation linear probing mIOU. For heterogeneous-resolution students, this is a seamless technique that allows ViT to operate at arbitrary resolutions within some envelope.
@@ 6 Conclusion and Key Insights
Dense features. As seen in figure 1, RADIO is capable of producing high resolution and low-noise features. Given the adaptor heads, it is able to reasonably reproduce the features generated by the teachers, and the backbone features themselves are highly coherent. An issue we identified, however, shown in figure 4 is that RADIO appears to have a latent ‘low resolution’ and ‘high resolution’ mode, likely due to the partitioned training between CLIP+DINO and SAM objectives, which we intend to fix in future work.
@@ Appendix F Visual Question Answering Samples
Figures 8 to 12 show sample questions from our Visual Question Answering datasets, together with sample answers when using our vision encoders in a LLaVA setup.

### v5

@@ Abstract
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 1: AM-RADIO is a framework to distill multiple pretrained vision foundation models, such as CLIP [51], DINOv2[48], SAM [35], into a single model that we call RADIO. As a result, a single vision foundation model agglomerates unique properties of the original models. This unifying approach obtains state-of-the-art feature representations in a single forward pass while also enabling unique properties such as zero-shot classification (CLIP) or open set instance segmentation (SAM) at negligible additional cost.
Image description: (left) PCA feature visualization of different models. Our proposed RADIO model can process any resolution and aspect ratio, and produces semantically rich dense encodings; (middle) the overview of the AM-RADIO framework; (right) benchmarks on classification, segmentation, and vision-language modeling tasks, see section 5.
@@ 4 Implementation Details
Multi-Resolution ViTs. Many of our student models use ViT [16] as the base vision architecture. Traditionally, ViTs use a learned position embedding for each input patch in an image, which in turn enforces that the model always operates at a constant resolution. We employ the Cropped Position Embedding (CPE) [33] augmentation with the number of positions being equal to $128^{2}$. The position embeddings are then randomly cropped and interpolated to match the number of input patches for the student model. Even when training with CLIP+DINOv2 at 224 resolution, we found that this technique results in a negligible drop (Table 5) in summary metrics, but improved semantic segmentation linear probing mIOU. For heterogeneous-resolution students, this is a seamless technique that allows ViT to operate at arbitrary resolutions within some envelope. In addition to enabling arbitrary resolutions, as shown in figure 3, CPE reduces the noise artifacts in the position embeddings as compared to other ViT models [67, 66, 6].
@@
(a) RADIO 2048px
(b) DINOv2-g-reg 518px
(c) DFN CLIP 378px
(d) OpenAI CLIP 336px
Figure 3: PCA visualization of the position embeddings for various models. The CPE method not only allows RADIO to learn an arbitrarily large absolution position embedding map, but also goes a long way towards regularizing the space and eliminating high frequency artifacts. As seen with the other models, position embeddings normally have regular frequency patterns, leading to undesirable output artifacts from the ViT [67, 66, 6].
@@ 5.4 Visual Question Answering
5.5 3D Awareness Probing
Following the work from [17], we probe our model’s ability to extract 3D features such as depth, surface normals and multi-view keypoint correspondance. Our results are summarized in Table 9 and show that our model’s performance is on par with the bigger DINOv2-g-14-reg[14] and significantly better than other comparably-sized teachers.
Backbone Depth Surface Multi-view
Normals corr.
DFN CLIP-H/14 52.5 23.0 20.3
OpenAI CLIP-L/14 53.7 25.3 20.7
DINOv2-g/14-reg 83.2 59.6 59.9
SAM-H/16 68.2 50.3 45.3
RADIO-ViT-H/16 (ours) 81.0 58.5 62.1
Table 9: Probing 3D Awareness: we use the code from [17] and evaluate our RADIO model and its teachers on monocular depth, surface normals and multi-view correspondance tasks, using the NAVI[31] dataset. For each task we report the accuracy, averaged over all thresholds.
@@ 6 Conclusion and Key Insights
Dense features. As seen in figure 1, RADIO is capable of producing high resolution and low-noise features. An issue we identified, however, shown in figure 5 is that RADIO appears to have a latent ‘low resolution’ and ‘high resolution’ mode, likely due to the partitioned training between CLIP+DINO and SAM objectives, which we intend to fix in future work.
@@ Appendix F Visual Question Answering Samples
Figures 9 to 13 show sample questions from our Visual Question Answering datasets, together with sample answers when using our vision encoders in a LLaVA setup.